# SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

### v1

@@ Abstract
Extreme low-bit quantization is critical for efficiently deploying Large Language Models (LLMs), yet it often leads to severe performance degradation at 2-bits and even 4-bits (e.g., MXFP4). We present SignRoundV2, a post-training quantization framework that is highly effective even without mixed-precision. SignRoundV2 introduces (1) a fast sensitivity metric that combines gradient information with quantization-induced deviations to guide layer-wise bit allocation, and (2) a lightweight pre-tuning search for quantization scales to improve extremely low-bit quantization. These components allow SignRoundV2 to close the gap with full-precision models. Extensive experiments indicate that our method sustains competitive accuracy for LLMs, achieving production-grade performance with $\sim$1% variance at $4\text{-}5$ bits and strong results even at 2 bits. The implementation is available at https://github.com/intel/auto-round.
SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs
Wenhua Chengthanks: Correspondence: wenhua.cheng@intel.com  and Weiwei Zhangthanks: These authors contributed equally  and Heng Guo22footnotemark: 2  and Haihao Shen Intel
@@ 1 Introduction
The advent of Large Language Models (LLMs) such as GPT (brown2020language), LLaMA (touvron2023llama; meta2024llama3), Qwen (bai2023qwen; yang2025qwen3) and DeepSeek (liu2024deepseek) has marked a major shift in modern artificial intelligence. As parameter counts scale from billions to hundreds of billions, these models have demonstrated unprecedented capabilities in reasoning, coding, multimodal understanding, and autonomous agent behaviors. However, such scaling comes at the cost of dramatically increased memory consumption, bandwidth pressure, and inference latency, posing significant barriers for real-world deployment, especially on resource-constrained environments such as consumer GPUs, CPUs, or edge devices.
@@
To alleviate these limitations, post-training quantization (PTQ) has emerged as one of the most practical approaches because it avoids costly re-training and can be applied to a wide range of pretrained LLMs (xiao2023smoothquant; frantar2022gptq; lin2024awq; cheng2024optimize). By compressing weights and activations to low-bit representations, PTQ can yield substantial reductions in memory footprint and hardware cost. In particular, the push toward extreme low-bit quantization, for example, sub-4-bit weight-only quantization and 4-bit activation quantization, has become a crucial enabler for democratizing LLM deployment. However, achieving this while preserving accuracy remains highly challenging due to increased quantization error, layer sensitivity imbalance, and the lack of hardware-friendly data types.
Figure 1: Average accuracy of pure 2-bit (W2A16) models on Llama 2/3 70B. See detailed results in Table 1.
@@
Recent research has moved beyond conventional W4A16 and W8A8 schemes (where WA denotes weight and activation bit-widths) toward more expressive low-bit data types. For example, MXFP4 (mxfp4) and NVFP4 (nvpf4) are newly introduced floating-point variants optimized for modern accelerators, demonstrating that well-designed data types can maintain reasonable accuracy even under aggressive activation quantization. In parallel, weight-only quantization continues to push the limits of bit efficiency. In particular, BitNet (wang2025bitnet; wu2025bitnet1) explores quantization down to approximately 1.58 bits, showing that extremely low-bit representations can still be learned if specialized training or distillation procedures are used. However, these methods require training from scratch or costly post-training distillation, making them less suitable for scenarios where fast, scalable, and hardware-efficient PTQ is desired.
@@
A natural direction for overcoming the limitations mentioned above is Adaptive Bit-Width Quantization (ABQ). ABQ mitigates the shortcomings of uniform-bit quantization by allocating higher precision to more sensitive layers while aggressively compressing more robust ones (zhou2018adaptive; dong2019hawq; wang2019haq; NEURIPS2024_2c30a37c). The central challenge therefore becomes the design of an accurate and efficient layer sensitivity metric to guide bit allocation. Early methods estimate sensitivity using the Hessian-related matrix (dong2019hawq; NEURIPS2024_2c30a37c), or rely on layer-wise/model-wise output distortion measured on a small calibration dataset (hubara2020improving). However, these heuristics often exhibit imperfect correlation when applied layer-wise, and model-wise approaches typically require slow calibration, especially for large language models (LLMs). In contrast, learning-based or reinforcement learning based approaches automate precision assignment (lou2019autoq; wang2019haq), but at the cost of substantial compute overhead, contradicting the primary objective of efficient PTQ.
@@
To address these challenges, we introduce SignRoundV2, an enhanced post-training quantization framework building upon the original SignRound (cheng2024optimize). Our method is driven by two key insights. First, an effective sensitivity metric must capture both the local parameter distortion caused by quantization and its global impact on task loss. We design a simple yet highly effective metric that integrates gradient information with quantization-induced parameter deviation, yielding a far more reliable signal for ABQ than existing proxies. Then given a target average bit budget, SignRoundV2 employs dynamic programming to convert layer-wise sensitivity scores into an optimal bit configuration. Second, we observe that the initialization of quantization parameters especially scale is critical for extremely low-bit stability, consistent with findings from recent PTQ literature (lin2024awq; ma2024era). Accordingly, inspired by the importance matrix in llama.cpp (ggerganov_llamacpp), we introduce a lightweight pre-tuning search that rapidly identifies high-quality initialization values before the main tuning stage, substantially improving recoverable accuracy. Figure 1 compares our method with state-of-the-art baselines, showing that it achieves similar performance at significantly lower cost and avoids the issues associated with Quantization Aware Training (QAT), which are discussed in Section 2. In summary, our contributions are three-fold.
  • We propose a simple and efficient sensitivity metric that integrates gradient information with quantization-induced parameter deviations. This metric captures both local parameter distortions and their global impact on task loss, providing a reliable guide for adaptive bit-width quantization.
  • We incorporate a lightweight pre-tuning search for quantization scales prior to the main tuning stage, enhancing both stability and final model accuracy under extremely low-bit settings. Experimental results demonstrate that this strategy achieves performance on par with recent efficient QAT methods.
  • We develop a fast, adaptive mixed-precision quantization method and show that it achieves competitive performance compared to existing QAT approaches, while incurring substantially lower quantization cost.
@@ 2 Related Work
QAT integrates learnable quantizers directly into the training loop, minimizing the task loss and fine-tuning additional parameters beyond the quantizers introduced by low-precision operations, as described in EfficientQAT (chen2025efficientqat), LLM-QAT (liu2023llm), DL-QAT (ke2024dl), and BitDistiller (du2024bitdistiller). Recent work(chee2023quip; tseng2024quip; egiazarian2024extreme) has revisited vector quantization for extremely low-bit settings, but these methods still rely on QAT to achieve high accuracy. Despite their effectiveness, QAT methods suffer from several practical limitations. First, they optimize the task loss or the unquantized parameters of the model, which can lead to catastrophic forgetting (luo2025empirical; kalajdzievski2024scaling), particularly in domains not covered by fine-tuning. Second, QAT requires careful hyperparameter tuning, and its actual quantization cost is often much higher than what is reported, typically far exceeding that of post-training quantization. Third, QAT generally requires substantially more data than PTQ to mitigate overfitting, further increasing the overall resource demands.
@@
Post-training quantization (PTQ) avoids additional fine-tuning of non-quantized parameters and task loss, and therefore offers a simple and resource-efficient pipeline for compressing LLMs (nagel2019data; liu2021post; frantar2022optimal; yao2021hawq; xiao2023smoothquant; wei2023outlier; dettmers2022gpt3). Because LLM decoding is typically memory-bound, weight-only PTQ has become the dominant approach (frantar2022gptq; lin2024awq; cheng2023teq; yao2024exploring; badri2023hqq; kim2023squeezellm; cheng2024optimize) in LLM inference. However, existing weight-only methods still suffer from a significant accuracy drop at extremely low bit-widths (e.g., 2-bit).
Mixed Precision Quantization.
Mixed-precision quantization assigns heterogeneous bit-widths across layers or channels to better align with their sensitivity. Early approaches include Hessian-based methods such as HAWQ (dong2019hawq) and BAQ (zhang2025baq), as well as reinforcement-learning–based schemes such as HAQ (wang2019haq). However, these methods are prohibitively expensive for multi-billion-parameter LLMs due to second-order matrix computations or extensive policy evaluations. More recent systems, such as MicroMix (liu2025micromix), apply FP4/FP6/FP8 at the channel level via MXFP microscaling kernels, but their reliance on specialized hardware limits the portability. In practice, Llama.cpp (ggerganov_llamacpp) adopts a more lightweight strategy that relies on heuristic rules to assign mixed precisions for each model and quantization scheme. Recent works (duanmu2025mxmoe; tao2025moqae; chitty2025mopeq) focus on mixed-precision settings for mixture-of-expert (MoE) models, which have become particularly popular for large-scale models.
@@
Rounding critically affects the quality of quantized weights. AdaRound (nagel2020up) formulates weight rounding as a per-weight binary optimization problem, where a second-order Taylor expansion of the task loss approximates quantization-induced perturbations, and a layer-wise local reconstruction loss is minimized via continuous relaxation. This formulation has inspired subsequent works such as BRECQ (li2021brecq), SignRound (cheng2024optimize). FlexRound (lee2023flexround) increases rounding flexibility through element-wise scaling, albeit at the cost of introducing numerous hyperparameters. Oscillation-Free training (liu2023oscillation) highlights instability arising from learnable rounding parameters. Activation-dependent rounding in AQuant (li2022efficient) reduces activation error but is not applicable to weight-only inference. SignRound (cheng2024optimize) jointly optimizes rounding and weight clipping using sign-based gradient descent, achieving substantial improvements, particularly at extremely low bit-widths such as 2 bits.
@@ 3 Methodology
This section presents our methodology. We begin in Section 3.1 by introducing the sensitivity metric, DeltaLoss, which guides the mixed-bit assignment. Section 3.2 then leverages DeltaLoss together with dynamic programming to generate adaptive bit-widths for each layer. In Section 3.3, we describe our pre-tuning search for quantization parameters. Finally, Section 3.4 lists all major hyperparameters and details a practical trick for loss calculation.
Before introducing our method, we first provide a brief overview of quantization and SignRoundV1  (cheng2024optimize). The following operation can be used to quantize and dequantize (qdq) the weights $\mathbf{W}$:
@@
SignRoundV1 (cheng2024optimize) introduces three trainable parameters $v$, $\alpha$ and $\beta$, to reduce the quantization error. The parameter $v$ enhances the rounding operation
@@
$s=\frac{\max(\mathbf{W})*\alpha-\min(\mathbf{W})*\beta}{2^{\text{bit}}-1}$ (4)
@@ 3.1 DeltaLoss Sensitivity Metric
Prior work  (dong2019hawq; NEURIPS2024_2c30a37c) largely ignores first-order gradient information and instead relies on second-order approximations, such as the Hessian or Fisher information matrix, to estimate sensitivity, under the assumption that gradients are typically near zero. However, this assumption breaks down when quantization induces a large loss. Moreover, although second-order metrics capture local curvature, they share a fundamental limitation with purely gradient-based methods: both fail to reflect the actual impact of quantization on the loss, providing only a relative trend rather than a reliable measure of true loss.
To address this limitation, we employ a first-order Taylor expansion to directly estimate the loss change induced by quantization, using the commonly adopted cross-entropy loss. Specifically, for a given layer with weight tensor $\mathbf{W}$ and activation tensor $\mathbf{A}$, the loss variation caused by quantizing the layer can be approximated using the first-order Taylor expansion:
$\Delta\mathcal{L}\approx g_{aq}\circ(A_{\text{f}}-A_{\text{q}})+g_{wq}\circ(W_{\text{f}}-W_{\text{q}})$ (5)
where:
  • $g_{aq}=\frac{\partial\mathcal{L}}{\partial A_{q}}$ represents the gradient of the loss with respect to the qdq activation tensor
  • $g_{wq}=\frac{\partial\mathcal{L}}{\partial W_{q}}$ represents the gradient of the loss with respect to the qdq weight tensor
  • $A_{\text{f}}$ and $W_{\text{f}}$ denote the full-precision activation and weight values, respectively
  • $A_{\text{q}}$ and $W_{\text{q}}$ denote the qdq activation and weight values, respectively
  • $\circ$ denotes the Hadamard product
In practice, we found that using the absolute value function yields more robust results, so we redefine the metric as
$\Delta\mathcal{L}\approx\left|(g_{aq}\circ(A_{\text{f}}-A_{\text{q}})\right|+\left|g_{wq}\circ(W_{\text{f}}-W_{\text{q}})\right|$ (6)
This formulation effectively combines gradient information and parameter distortion to yield a more accurate and holistic sensitivity estimate for each layer.
To reduce memory consumption, activation quantization has been identified as the dominant contributor to quantization loss as shown in (xiao2023smoothquant; wei2023outlier). Therefore, within the context of joint weight–activation quantization, we primarily focus on activation-induced distortion and omit the weight quantization term in our sensitivity formulation.
$\Delta\mathcal{L}$ $\approx\left|g_{aq}\circ(A_{\text{f}}-A_{\text{q}})\right|$ (7)
@@
Figure 2 visualizes the layer-wise sensitivity measured by DeltaLoss. It is evident that sensitivity varies across different data types, making it prohibitively labor-intensive to rely on heuristic rules to manually set quantization parameters for each scheme, as done in llama.cpp (ggerganov_llamacpp). Moreover, certain layers within a block, such as down_proj, exhibit higher sensitivity, suggesting that traditional block-wise fallback strategies may be suboptimal.

### v2

# SignRoundV2: Toward Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

@@ Abstract
Extremely low-bit quantization is critical for efficiently deploying Large Language Models (LLMs), yet it often leads to severe performance degradation at 2 bits and even at 4 bits (e.g., MXFP4). We present SignRoundV2, a post-training quantization framework designed to maintain high performance even under aggressive compression. SignRoundV2 introduces (1) a simple yet efficient adaptive mixed-precision strategy that leverages gradient information and quantization-induced reconstruction errors to guide layer-wise bit allocation, and (2) a set of lightweight stabilization techniques, including loss filtering and a pre-tuning scale search, to improve tuning effectiveness in extremely low-bit regimes. Our approach takes a significant step toward closing the performance gap between quantized and full-precision models. Experimental results across diverse LLMs demonstrate that SignRoundV2 achieves near-lossless performance in mixed MXFP settings, narrowing the gap to $\sim$1% at an average of 4.5 bits, while substantially improving accuracy in challenging 2-bit weight-only quantization. The source code is available at https://github.com/intel/auto-round.
11footnotetext: These authors contributed equally.22footnotetext: Correspondence: wenhua.cheng@intel.com33footnotetext: Working done during the internship at Intel.
@@ 1 Introduction
The advent of Large Language Models (LLMs) such as GPT [8], LLaMA [60, 62], Qwen [4, 70] and DeepSeek [41] has marked a major shift in modern artificial intelligence. As parameter counts scale from billions to hundreds of billions, these models have demonstrated unprecedented capabilities in reasoning, coding, multimodal understanding, and autonomous agent behaviors. However, such scaling comes at the cost of dramatically increased memory consumption, bandwidth pressure, and inference latency, posing significant barriers for real-world deployment, especially in resource-constrained environments such as consumer GPUs, CPUs, and edge devices.
@@
To alleviate these limitations, post-training quantization (PTQ) [69, 22, 39, 12] has emerged as one of the most practical approaches because it avoids costly re-training and can be applied to a wide range of pretrained LLMs. By compressing weights and activations to low-bit representations, PTQ yields substantial reductions in memory footprint and hardware cost. In particular, the push toward extremely low-bit quantization, for example, sub-4-bit weight-only quantization and 4-bit activation quantization, has become a crucial enabler for democratizing LLM deployment.
(a) W2A16 results on Llama 2/3 70B
(b) Recovery ratio at 2.5 bits
Figure 1: Left: results on pure 2 bits (W2A16) for Llama 2/3 70B. Right: recovery ratio for adaptive bit assignments at 2.5 bits across four models.
@@
Recent research has moved beyond conventional W4A16 and W8A8 schemes (where W and A denote weight and activation bit-widths) toward more expressive low-bit data types. For example, MXFP4 [1] and NVFP4 [54] are floating-point variants provided by modern accelerators, demonstrating that well-designed data types can maintain reasonable accuracy even under aggressive activation quantization. Despite these architectural advances, achieving high-accuracy recovery remains non-trivial, primarily due to the prevalent outliers in large-scale model. In parallel, weight-only quantization continues to push the limits of bit efficiency. In particular, BitNet [64, 68] explores quantization down to approximately 1.58 bits, showing that extremely low-bit representations can still be learned with specialized training or distillation. However, these methods require training from scratch or costly post-training distillation, making them less suitable for scenarios where fast, scalable, and hardware-efficient PTQ is desired.
@@
To minimize the accuracy loss caused by quantization, adaptive mixed-precision quantization has emerged as a natural solution to the limitations of uniform quantization. By strategically allocating higher precision to sensitive layers while aggressively compressing more robust ones, it effectively balances model performance with computational efficiency [77, 17, 65, 73]. Therefore, the central challenge is the design of an accurate and efficient layer sensitivity metric to guide bit allocation. Early methods estimate sensitivity using Hessian-based metrics [17, 73, 53], or rely on layer-wise or model-wise output distortion measured on a small calibration dataset [30]. However, these heuristics often exhibit imperfect correlation when applied layer-wise, and model-wise approaches typically require slow calibration, especially for large language models (LLMs). In contrast, learning-based or reinforcement-learning-based approaches automate precision assignment [47, 65, 35], but at the cost of substantial compute overhead.
@@
To address these challenges, we introduce SignRoundV2, an enhanced post-training quantization framework building upon SignRound [12]. Our method first utilizes a simple yet efficient adaptive mixed-precision strategy that integrates gradient information with quantization-induced deviations to capture both local parameter distortion and global task impact. By leveraging these reliable sensitivity signals, SignRoundV2 employs dynamic programming [5] to efficiently determine the optimal layer-wise bit configuration for a given budget, substantially narrowing the performance gap between quantized and full-precision models even under aggressive compression. Second, we emphasize that stabilizing the optimization process is important in extremely low-bit regimes. Building on findings that quantization parameter initialization is critical [39, 49], we introduce a lightweight pre-tuning search, inspired by the importance matrix in llama.cpp [25], to rapidly identify high-quality initial scales before the main tuning stage. To further ensure optimization stability, we incorporate a loss filtering mechanism that excludes loss outliers during tuning, preventing outlier gradients from degrading the learned quantization parameters. Figure 1 compares our method to state-of-the-art baselines, showing that it achieves similar performance at significantly lower cost and avoids the issues associated with Quantization-Aware Training (QAT), which are discussed in Section 2. In summary, our contributions are three-fold.
  • We introduce an efficient adaptive mixed-precision strategy that leverages gradient information and quantization-induced deviations to guide layer-wise bit allocation via dynamic programming, maintaining high performance even under aggressive compression.
  • We employ simple yet effective stabilization techniques, such as loss filtering and a lightweight pre-tuning scale search, to improve stability and accuracy in extremely low-bit regimes.
  • We demonstrate that SignRoundV2 achieves competitive accuracy compared with recent high-cost QAT and vector-quantization methods in weight-only quantization, while requiring substantially lower quantization cost and maintaining strong performance in weight-activation scenarios.
@@ 2 Related Work
QAT integrates learnable quantizers directly into the training loop, minimizing the task loss and fine-tuning additional parameters beyond the quantizers introduced by low-precision operations, as described in EfficientQAT [10], LLM-QAT [44], DL-QAT [32], and BitDistiller [18]. Recent works [9, 63, 20] have revisited vector quantization for extremely low-bit settings, but these methods still rely on QAT to achieve high accuracy. Despite their effectiveness, QAT methods suffer from several practical limitations. First, because they optimize the task loss and may update non-quantizer parameters, they can be more sensitive to the fine-tuning data distribution and may introduce forgetting or domain-specific overfitting [48, 31]. Second, QAT often requires careful hyperparameter tuning, and the practical quantization cost can be substantially higher than that of PTQ. Third, QAT typically relies on more training data than PTQ to achieve stable generalization, further increasing the overall resource requirements.
@@
Post-training quantization (PTQ) avoids task-loss optimization and additional fine-tuning of model parameters and therefore offers a simple and resource-efficient pipeline for compressing LLMs [52, 46, 21, 71, 69, 66, 16]. Because LLM decoding is typically memory-bound, weight-only PTQ has become the dominant approach [22, 39, 11, 72, 3, 33, 12] in LLM inference. However, existing weight-only methods still suffer from a significant accuracy drop at extremely low bit-widths (e.g., 2-bit). Meanwhile, joint weight–activation quantization has emerged as a complementary direction, with techniques such as rotation-based transformations [9, 2, 45, 38, 58] proposed to address activation outliers, albeit at the cost of additional inference overhead.
Adaptive Mixed-Precision Quantization.
Adaptive mixed-precision quantization assigns heterogeneous bit-widths across layers or channels to better align with their varying sensitivity to quantization error. Early frameworks utilized second-order information, such as Hessian-based methods (e.g., HAWQ [17] and BAQ [75]), or reinforcement learning schemes like HAQ [65]. However, these approaches are prohibitively expensive for multi-billion-parameter LLMs due to the computational intensity of second-order matrix computations or extensive policy evaluations. While MixLLM [76] attempts to combine first- and second-order metrics for channel-level precision, its reliance on naive thresholding and neglect of activation distortion in weight-activation joint quantization often lead to suboptimal performance. To address these scalability issues, recent advances focus on optimizing the search space or eliminating search altogether. AMQ [35] treats mixed-precision assignment as an AutoML problem, employing search space pruning and quantization proxies to accelerate the process; nonetheless, it remains relatively slow for large-scale models. In contrast, SFMP [53] completely bypasses the computationally expensive search process by utilizing a fine-grained, hardware-friendly assignment logic. Similarly, MicroMix [43] applies FP4/FP6/FP8 precision at the channel level via MXFP microscaling kernels. While these methods achieve high efficiency, they often require specific hardware or specialized kernel support. In practical deployment, frameworks like llama.cpp [25] adopt lightweight heuristic strategies to assign mixed precision across different model architectures. Furthermore, given the rising prominence of Mixture-of-Experts (MoE) architectures, recent works such as MxMoE [19], MoQAE [59], and MoPEQ [13] have extended adaptive mixed-bit settings specifically to MoE models.
@@
Rounding critically affects the quality of quantized weights. AdaRound [51] formulates weight rounding as a per-weight binary optimization problem, where a second-order Taylor expansion of the task loss approximates quantization-induced perturbations, and a layer-wise local reconstruction loss is minimized via continuous relaxation. This formulation has inspired subsequent works such as BRECQ [36] and SignRound [12]. FlexRound [34] increases rounding flexibility through element-wise scaling, albeit at the cost of introducing numerous hyperparameters. Oscillation-Free training [42] highlights instability arising from learnable rounding parameters. Activation-dependent rounding in AQuant [37] reduces activation error but is not applicable to weight-only inference. SignRound [12] jointly optimizes rounding and weight clipping using sign-based gradient descent, achieving substantial improvements, particularly at extremely low bit-widths such as 2 bits.
@@ 3 Methodology
This section presents our methodology. We begin in Section 3.1 by introducing the sensitivity metric, DeltaLoss, which guides mixed-bit assignment. Section 3.2 then leverages DeltaLoss with dynamic programming to determine adaptive bit-widths for each layer. Section 3.3 describes the tuning strategy for quantization parameters, while Section 3.4 introduces our loss exclusion strategy. Finally, Section 3.5 summarizes the main hyperparameters and presents a practical trick for loss computation.
Before introducing our method, we briefly review the standard quantize-dequantize (QDQ) formulation and SignRoundV1 [12]. The QDQ operator for weights $\mathbf{W}$ is defined as follows:
@@
SignRoundV1 [12] introduces three trainable parameters $v$, $\alpha$ and $\beta$, to reduce the quantization error. The parameter $v$ enhances the rounding operation
@@
$s=\frac{\max(\mathbf{W})\cdot\alpha-\min(\mathbf{W})\cdot\beta}{2^{\text{bit}}-1}.$ (4)
@@ 3.1 DeltaLoss Sensitivity Metric
Prior work [17, 73] largely ignores first-order gradient information and instead relies on second-order approximations, such as the Hessian or Fisher information matrix, to estimate sensitivity under the assumption that gradients are near zero at the local optimum. However, this assumption breaks down when quantization induces a large loss. Moreover, although second-order metrics capture local curvature, they share a fundamental limitation with purely gradient-based methods: both fail to reflect the actual impact of quantization on the loss and therefore provide only a relative trend rather than a reliable measure of the true loss.
To address this limitation, we estimate the loss increase induced by the QDQ operation using a first-order Taylor approximation. Let $\mathcal{L}_{\mathrm{f}}$ and $\mathcal{L}_{\mathrm{q}}$ denote the losses of the full-precision and quantized models, respectively. We define the quantization-induced loss increase as
$\Delta\mathcal{L}=\mathcal{L}_{\mathrm{q}}-\mathcal{L}_{\mathrm{f}}.$ (5)
For a layer with full-precision weight tensor $W_{\mathrm{f}}$ and activation tensor $A_{\mathrm{f}}$, and their quantized counterparts $W_{\mathrm{q}}$ and $A_{\mathrm{q}}$, we approximate the loss increase by a first-order Taylor expansion around the quantized point:
$\Delta\mathcal{L}\approx g_{aq}^{\top}(A_{\mathrm{q}}-A_{\mathrm{f}})+g_{wq}^{\top}(W_{\mathrm{q}}-W_{\mathrm{f}}),$ (6)
where the gradients $g_{aq}=\partial\mathcal{L}/\partial A_{\mathrm{q}}$ and $g_{wq}=\partial\mathcal{L}/\partial W_{\mathrm{q}}$ are computed using the quantized model on a small calibration set.
While Eq. 6 provides a signed first-order estimate, positive and negative element-wise contributions may cancel each other out when aggregated over tokens, channels, and calibration samples. We therefore instead use the magnitude of the gradient-weighted perturbation as a more stable sensitivity score:
$\mathrm{DeltaLoss}=\left\|g_{aq}\odot(A_{\mathrm{q}}-A_{\mathrm{f}})\right\|_{1}+\left\|g_{wq}\odot(W_{\mathrm{q}}-W_{\mathrm{f}})\right\|_{1},$ (7)
where $\odot$ denotes element-wise multiplication and $\|\cdot\|_{1}$ denotes summation over all elements. The final layer-wise score is averaged over calibration samples.
This metric combines two complementary factors: the magnitude of the quantization perturbation and the task-aware gradient sensitivity. Layers that exhibit both large quantization deviations and large gradients receive higher DeltaLoss scores and are therefore assigned higher precision under a fixed bit budget.
For joint weight-activation quantization, both weight and activation perturbations can be included as in Eq. 7. In practice, prior studies [69, 66] indicate that activation quantization often dominates the degradation. Therefore, for memory-efficient mixed weight-activation settings where weights and activations share the same bit-width, we use the activation-induced term as the default sensitivity score:
$\mathrm{DeltaLoss}^{\mathrm{WA}}=\left\|g_{aq}\odot(A_{\mathrm{q}}-A_{\mathrm{f}})\right\|_{1}.$ (8)
@@
Figure 2 visualizes the layer-wise sensitivity measured by DeltaLoss. It is evident that sensitivity varies across different data types, making it prohibitively labor-intensive or suboptimal to rely on heuristic rules to manually assign quantization bit-widths for each scheme. Moreover, certain layers within a block, such as down_proj, exhibit consistently higher sensitivity, suggesting that traditional block-wise fallback strategies may be suboptimal.